OpenStudy (anonymous):

(Help?) Choose the point-slope form of the equation below that represents the line that passes through the points (-3, 2) and (2, 1). y + 3 = -5(x - 2) y - 2 = -5(x + 3) y + 3 = -one fifth(x - 2) y - 2 = -one fifth(x + 3)

OpenStudy (anonymous):

Point-slope form is (y-y1)=m(x-x1) First you would have to find the slope between those points. Use the slope formula y2-y1/x2-x1 (1-2)/ (2-(-3)) -1/5= the slope or m. We can cross the first 2 out because they have a slope of -5. Your answer would be y+3= -1/5(x-2) because that is the only pair that fits.

OpenStudy (anonymous):

ohhh okay! I get it thanks so its the 3rd choice ?

OpenStudy (anonymous):

@WordGEEK

OpenStudy (anonymous):

Yes it would be the third choice.